OpenStudy (anonymous):

the radius of a circle is 2π/300. Find the length of the arc that is formed by an angle of π/6 subtended at the centre of the circle

OpenStudy (anonymous):

so radius = pi/150. And angle = Pi/6 radians which on converting to degrees = 180/6 = 30. So theta(degrees) = 30, radius = Pi/150, Length of arc = theta x radius x Pi / 180

OpenStudy (anonymous):

thats exactly what i did but the answer is wrong :/ the question actually doesnt give a radius, they give the circumference which is 300mm have i calculated the radius incorrectly?

OpenStudy (anonymous):

if Circumference = 300mm, then 2 x Pi x r = 300; Pi x r = 150. R = 150/Pi. Length of arc = R X Theta x Pi/180 => 150x30xPi/Pix180 => 4500/180 = 25mm.

OpenStudy (anonymous):

thats the answer thanks!